### v1

# Uni3C: Unifying Precisely 3D-Enhanced Camera and Human Motion Controls for Video Generation

Abstract
Camera and human motion controls have been extensively studied for video generation, but existing approaches typically address them separately, suffering from limited data with high-quality annotations for both aspects. To overcome this, we present Uni3C, a unified 3D-enhanced framework for precise control of both camera and human motion in video generation. Uni3C includes two key contributions. First, we propose a plug-and-play control module trained with a frozen video generative backbone, PCDController, which utilizes unprojected point clouds from monocular depth to achieve accurate camera control. By leveraging the strong 3D priors of point clouds and the powerful capacities of video foundational models, PCDController shows impressive generalization, performing well regardless of whether the inference backbone is frozen or fine-tuned. This flexibility enables different modules of Uni3C to be trained in specific domains, i.e., either camera control or human motion control, reducing the dependency on jointly annotated data. Second, we propose a jointly aligned 3D world guidance for the inference phase that seamlessly integrates both scenic point clouds and SMPL-X characters to unify the control signals for camera and human motion, respectively. Extensive experiments confirm that PCDController enjoys strong robustness in driving camera motion for fine-tuned backbones of video generation. Uni3C substantially outperforms competitors in both camera controllability and human motion quality. Additionally, we collect tailored validation sets featuring challenging camera movements and human actions to validate the effectiveness of our method.
[…]
1 Introduction
Figure 1: Given a single-view image across various domains (e.g., real-world, text-to-image, animation), we first extract the monocular depth and focal length of it via Depth-Pro [6] and then achieve relative point clouds. Then, the proposed Uni3C can generate impressive videos under arbitrary camera trajectories (a), human motion characters (SMPL-X [36]), or both of these conditions (b). (c) Uni3C further supports the camera-controlled motion transfer.
Recent advancements in foundational video diffusion models (VDMs) [5, 64, 7, 27, 46, 26, 54] have unlocked unprecedented capabilities in creating dynamic and realistic video content. A significant challenge in this field is achieving controllable video generation, a feature with broad applications in virtual reality, film production, and interactive media. In this paper, we focus on two aspects of controllable video generation: camera control [17, 63, 69, 2, 66, 44] and human motion control [21, 57, 62, 72, 51, 22]—both of which are critical and interdependent in real-world scenarios.
Recent pioneering works have extensively studied controlling camera trajectories for VDMs through explicit conditions like Plücker ray [17, 2, 69, 18, 30], point clouds [66, 44, 28, 12, 40]. Concurrently, controllable human animation also attracted a lot of attention based on poses [21, 62, 51, 22] or SMPL characters [72, 75, 73]. Despite these advancements, several challenges remain: 1) Most approaches deeply hack the inherent capacities of VDMs, which have been trained with domain-specific data and conditions, inevitably undermining the generalization to handle out-of-distribution scenarios. 2) Very few works investigate the joint control of both camera trajectories and human motions. This requires diverse camera trajectories in human-centric videos with high-quality annotations [57], which are often expensive to obtain. 3) There is a lack of explicit and synchronized guidance that incorporates strong 3D-informed priors to concurrently control both camera movements and human motions. Relying on separate conditions, like point clouds and SMPL, struggles to represent physically reasonable interactions and positional relations between characters and environments.
To address these challenges, we present Uni3C, a novel framework that Unifies precise 3D-enhanced camera and human motion Controls for video generation as shown in Figure 1 via two key innovations. Firstly, we propose the insight that controlling camera trajectories for powerful foundational VDMs can be achieved by lightweight, trainable modules with rich informed guidance and reasonable training strategies. By avoiding hacking the underlying capacities of VDMs, our camera-controlling model can be directly generalized to versatile downstream tasks rather than costly joint training and extensive Structure from Motion (SfM) labeling [47, 68, 29]. Secondly, we claimed that camera and human motion controls are inherently interdependent. Therefore, we propose to align their conditions into a global 3D world during the inference phase, enabling 3D consistent generation across both domains.
Formally, Uni3C is built upon the foundational VDM—Wan2.1 [54]. For the control of camera trajectories, we propose PCDController, a plug-and-play control module with only 0.95B parameters (compared to 12B of Wan2.1) that operates on unprojected 3D point clouds derived from monocular depth estimation [6]. Thanks to the rich geometric priors of point clouds, PCDController is capable of fine-grained camera control, even when trained on constrained multi-view images and videos with a frozen backbone. Furthermore, PCDController can be compatible with fine-tuned VDM backbones for versatile downstream tasks. This surprising factor supports domain-specific training, i.e., camera and human motion modules can be trained independently without jointly annotated data. For the global 3D world guidance, we align scenic point clouds (for camera control) and SMPL-X characters [36] (for human animation) into the same world-coordinate space via the rigid transformation [52], while the 2D keypoints [61] bridge the relation of two presentations. Note that our alignment enables complicated motion transfer, covering disparate characters, positions, and viewpoints as verified in the last row of Figure 1.
Extensive experiments validate the efficacy of Uni3C. To evaluate the remarkable generalization of PCDController, we collect an out-of-distribution test set across different domains, where each image has four different camera trajectories. For the joint controllability, we build a comprehensive test set of in-the-wild human videos. GVHMR [48] is used to extract SMPL-X as the condition, while three complex and random camera trajectories are assigned for each video. VBench++ [23] is employed to verify the overall performance of our method. Uni3C significantly outperforms other competitors, both quantitatively and qualitatively.
We highlight the key contributions of Uni3C as:
  • PCDController. A robust, lightweight camera control module is proposed, which enjoys strong 3D priors from point clouds, compatible with both frozen or adapted VDMs.
  • Global 3D World Guidance. A unified inference framework that aligns scene geometry (point clouds) and human characters (SMPL-X) for 3D-coherent video control.
  • Comprehensive Validation. We propose new benchmarks and datasets to evaluate challenging camera-human interaction scenarios, demonstrating Uni3C’s superiority over existing approaches.
2 Related Work
Camera Control for VDMs.
Controlling camera trajectories in video generation has garnered significant attention in recent times. Some works focused on injecting camera parameters into VDMs to achieve camera controllability [58, 17, 3, 2, 69, 18, 30], typically utilizing the Plücker ray presentation. For instance, VD3D [3] designed a tailored framework for Diffusion Transformer (DiT) [38], while AC3D [2] further emphasized the generalization with less trainable parameters. Moreover, DimensionX [50] further decoupled the spatial and temporal control with different LoRAs [20]. Despite the progress made by these methods, they cannot control the camera movements precisely, particularly when the case is beyond the training domains with an unknown metric scale. Thus, other recent works have employed point cloud conditions through both training-based [66, 44, 28, 12, 40] and training-free [19, 65] manners. However, these methods fail to accommodate the generalized model design and training strategy to handle imperfect point clouds or out-of-distribution data, especially in scenarios involving humans or animals.
Unified Control for VDMs.
Recent works have unified multiple conditions to guide video generation [58, 12, 13, 57, 10, 14, 70]. MotionCtrl [58] integrated the camera and object motion controls through separate pose and trajectory injections. Subsequently, researchers further explored the presentation of conditions, such as point trajectories [12], point tracking [13, 14], and 3D-aware signals [10]. Humanvid [57] first unified the control of human scenes, while VidCraft3 [70] further considered the lighting control. Although these pioneering approaches achieved promising performance, they often rely heavily on joint training with various conditions and well-labeled datasets. Furthermore, there has been limited discussion on unified control within foundational VDMs that exceed 10B parameters. Our work offers a solution to address these issues: unifying existing models for different downstream tasks without the need for costly fine-tuning or fully annotated datasets. This strategy is particularly well-suited for large VDMs, enabling models to focus more on enhancing performance within their specific domains.
3 Preliminary: Video Diffusion Models
We briefly review VDMs and Wan2.1 [54] in this section as preliminary knowledge. VDMs are mainly based on the latent diffusion model [45], modeling the conditional distribution $p(z_{0}|c_{txt},c_{img})$, where $z_{0}$ indicate clean video latent features encoded by 3D-VAE; $c_{txt},c_{img}$ denote the text condition for Text-to-Video (T2V) and the optional image condition for Image-to-Video (I2V), respectively. The training of VDM involves reversing the diffusion process by the noise estimator $\epsilon_{\theta}$ as:
$\min_{\theta}\mathbb{E}_{z_{0},t,\epsilon,c_{txt},c_{img}}[\|\epsilon_{\theta}(z_{t},t,c_{txt},c_{img})-\epsilon\|^{2}],$ (1)
where $\epsilon\sim\mathcal{N}(0,I)$ indicates Gaussian noise; timestep $t\in[0,T_{max}]$; $z_{t}$ is the intermediate noisy latent state of timestep $t$. Recently, most VDMs have employed Flow Matching (FM) [33] as the improved diffusion process with faster convergence and more stable training. Based on the ordinary differential equations (ODEs), FM formulates the linear interpolation between $z_{0}$ and $z_{1}$, i.e., $z_{t}=tz_{1}+(1-t)z_{0}$, where $t\in[0,1]$ is sampled from the logit-normal distribution. The velocity prediction $v_{\theta}$ can be written as:
$\min_{\theta}\mathbb{E}_{z_{0},t,\epsilon,c_{txt},c_{img}}[\|v_{\theta}(z_{t},t,c_{txt},c_{img})-v_{t}\|^{2}],$ (2)
where the ground truth velocity denotes $v_{t}=\frac{dz_{t}}{dt}=z_{1}-z_{0}$. Additionally, recent foundational VDMs [7, 64, 26, 54] are built with DiT [38] to achieve more capacities for video generation.
Wan2.1 [54] is an open-released, large-scale VDM with DiT architecture trained with flow matching [33]. umT5 [11] is utilized as the multi-language text encoder to inject textual features into Wan2.1 through cross-attention. For image-to-video, Wan-I2V further incorporates features from CLIP’s image encoder [42] to improve the results. Uni3C is primarily designed for Wan-I2V with 14B parameters, but we empirically find that it is compatible with the Wan-T2V version as verified in Section 5.4, showing convinced generalization of PCDController.
4 Method
Figure 2: The overview pipeline of PCDController. PCDController is built as a lightweight DiT trained from scratch. We first obtain the point clouds via the monocular depth extracted from the first view. Then, the point clouds are warped and rendered into the video $V_{pcd}$. The input conditions for PCDController comprise rendered $V_{pcd}$, Plücker ray $\mathbf{P}$, and the noisy latent $z_{t}$. Note that only the PCDController and camera encoder are trainable in our framework.
Overview.
Given a reference view $I_{img}\in\mathbb{R}^{3\times h\times w}$, camera trajectories $\{c_{cam}^{i}\}_{i=1}^{N}$ of $N$ target views, and textual condition, Uni3C uses PCDController to produce the target video $\{V_{tar}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 3\times h\times w}$ under specified camera trajectories. This can be formulated as $p(z_{0}|c_{txt},c_{img},c_{cam})$, where $z_{0}$ indicates the clean latent video feature, $c_{txt},c_{img}$ are textual features from umT5 and image latent condition encoded from $I_{img}$, respectively. We show the overview pipeline of PCDController in Figure 2, which is a core component of Uni3C for generalized camera control (Section 4.1) across diverse downstream tasks. In this work, we focus on human animation (Section 4.2). Subsequently, we introduce the global 3D world guidance illustrated in Figure 5 to unify both camera and human characters into a consistent 3D space for inference (Section 4.3).
4.1 PCDController with 3D Geometric Priors
Architecture.
Following AC3D [2], PCDController is designed within a simplified DiT module rather than copying modules and weights from the main backbone as ControlNet [67]. To preserve the generalization of Wan-I2V, we follow the insight of training as few parameters as possible once the effective camera control has been achieved. Formally, we reduce the hidden size of PCDController from 5120 to 1024, while zero-initialized linear layers are used to project hidden size back to 5120 before being added to Wan-I2V. Moreover, as investigated in [2, 30], VDMs mainly determine camera information through shallow layers. Thus, we only inject camera-controlling features into the first 20 layers of Wan-I2V to further simplify the model. Additionally, we discard the textual condition for PCDController to alleviate intractable hallucination and remove all cross-attention modules. In this way, the overall number of trainable parameters for PCDController is reduced to 0.95B, a significant reduction compared to Wan-I2V (14B).
3D Geometric Priors.
In contrast to merely utilizing Plücker ray as the camera embedding [2, 30], we incorporate much stronger 3D geometric priors to compensate the model simplification, i.e., videos $\{V_{pcd}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 3\times h\times w}$ rendered from point clouds under given camera trajectories. Specifically, we first use Depth-Pro [6] to extract the monocular depth map from the reference view. We then align this depth map into a metric representation using SfM annotations [47] or multi-view stereo [8]. Following [9], we employ RANSAC to derive the rescale and shift coefficients, preventing the collapse of constant depth outcomes. Subsequently, the point clouds $X_{pcd}\in\mathbb{R}^{hw\times 3}$ are got by unprojecting all 2D pixels from $I_{img}$ into the world coordinate via its metric depth $\hat{D}_{img}$ as follows:
$X_{pcd}(x)\simeq R_{c\rightarrow w}\hat{D}_{img}(x)K^{-1}\hat{x},$ (3)
where $x$ denotes the 2D coordinates of $I_{img}$, while $\hat{x}$ refers to the homogeneous coordinates; $K,R_{c\rightarrow w}$ mean the intrinsic and extrinsic cameras of the reference view, respectively. After that, we render $\{V_{pcd}^{i}\}_{i=2}^{N}$ for the remaining ($N-1$) views by PyTorch3D according to their respective camera intrinsics and extrinsics. Note that the first rendering corresponds to the reference image, i.e., $V_{pcd}^{1}=I_{img}$, to confirm the identity. We apply $V_{pcd}$ to the 3D-VAE of Wan2.1 to achieve $c_{pcd}$ as the point latent condition. To further handle the significant viewpoint changes, which may extend beyond the point clouds’ visibility of the first frame, PCDController also includes Plücker ray embedding [60], $\{\mathbf{P}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 6\times h\times w}$, as the auxiliary condition. $\{\mathbf{P}^{i}\}_{i=1}^{N}$ is encoded by a small camera encoder, comprising casual convolutions and a 4-8-8 downsampling factor to keep the same sequential length as 3D-VAE outputs. Therefore, the distribution modeling of PCDController can be written as $p(z_{0}|c_{txt},c_{img},c_{pcd},\mathbf{P})$.
Figure 3: Results of PCDController with imperfect point clouds. Benefiting from the well-preserved capacity from VDM, PCDController enjoys robust generation with inferior point clouds.
Discussion.
We empirically find that our method demonstrates robust performance even when working with imperfect point clouds obtained through monocular depth unprojection. In this context, the point clouds serve as the primary camera control signal, facilitating the convergence of training rather than dominating the multi-view geometric and textural generations, as illustrated in Figure 3. By achieving earlier model convergence, we can better maintain the inherent capabilities of VDMs.
4.2 Human Animation
In this paper, we explore the unified control through two human animation approaches, both of which are built on the Wan2.1 framework, targeting I2V and T2V, respectively. While these methods are not the primary focus of our work, we provide a brief introduction here. Formally, the concurrently pioneering work, RealisDance-DiT [73], is used to replace the Wan-I2V backbone for high-quality human animation. RealisDance-DiT incorporates SMPL-X [36] and Hamer [37] as additional input conditions to guide human motions. To ensure the flexibility for motion transfer, RealisDance-DiT randomly selects the reference frame in the video sequence, which is not perfectly aligned with the given SMPL-X. RealisDance-DiT only trains self-attention modules and patchify encoders to confirm the generalization. However, we clarify that integrating the control branch trained within different backbones is still challenging, which is addressed by our elaborate PCDController. Furthermore, we tried another version, RealisDance-DiT-T2V, based on Wan-T2V without reference image conditions to explore the generalization of PCDController. Remarkably, PCDController adapts successfully to Wan-T2V, empowering it with impressive I2V ability.
4.3 Global 3D World Guidance
Figure 4: The assignment of multi-modal conditions for Uni3C. Camera, point clouds, SMPL-X [36] and Hamer [37] are all presented as 3D conditions.
Figure 5: The overview of global 3D world guidance. (a) We first align the SMPL-X characters from the human world space $W_{hum}$ to the environment world space $W_{env}$ comprising dense point clouds. (b) GeoCalib [53] is used to calibrate the gravity direction of SMPL-X. (c) The rigid transformation coefficients $\tilde{s},\tilde{R},\tilde{t}$ are employed to align the whole SMPL-X sequence. We re-render all aligned conditions under specific camera trajectories as the global 3D world guidance.
Definition.
As illustrated in Figure 4, Uni3C adopts multi-modal conditions, including camera, point clouds, reference image, SMPL-X [36], and Hamer [37]. The first three conditions are used for PCDController, while the latter three are used for human animations. We employ GVHMR [48] to recover SMPL-X characters. One can also retrieve desired motion sequences from motion datasets [39, 41, 16, 31] or generate new motions through text-to-motion models [24, 15]. Although most of the conditions above are formulated as 3D presentations, they stay in two different world coordinates. We define the point cloud world coordinate as the environmental world space $W_{env}$, which is the main world space controlled by cameras, while the SMPL-X is placed in the human world space $W_{hum}$. It is non-trivial to control the camera across two different world spaces consistently. For example, determining the initial camera position within $W_{hum}$ is particularly ambiguous, especially for tasks involving motion transfer. Therefore, we propose the global 3D world guidance that places the human condition into the “environment”, i.e., aligning SMPL-X from $W_{hum}$ to $W_{env}$ as shown in Figure 5(b).
Multi-Modal Alignment.
Fortunately, 2D human pose keypoints subtly bridge $W_{hum}$ and $W_{env}$. Formally, we first estimate 17 human keypoints $\{\mathbf{k}^{i}_{2D}\}_{i=1}^{17}\in\mathbb{R}^{17\times 2}$ from the reference view $I_{img}$ by ViTPose++ [61]. Then, we unproject 2D keypoints into $W_{env}$ to obtain 3D keypoints $\{\mathbf{k}^{i}_{env}\}_{i=1}^{17}\in\mathbb{R}^{17\times 3}$ through metric monocular depth $\hat{D}_{img}$ and the intrinsic camera of the reference image. For SMPL-X in $W_{hum}$, the COCO17 regressor [48] is utilized to project the first SMPL-X character into $\{\mathbf{k}^{i}_{hum}\}_{i=1}^{17}$ corresponding to the same human keypoints. Consequently, a least-squares estimation [52] based rigid transformation can be used to align $\mathbf{k}_{hum}$ to $\mathbf{k}_{env}$ as follows:
$\min_{\tilde{s},\tilde{R},\tilde{t}}\sum_{i=1}^{17}w_{i}\|(\tilde{s}\tilde{R}(\mathbf{k}^{i}_{hum})^{T}+\tilde{t})^{T}-\mathbf{k}^{i}_{env}\|^{2},$ (4)
where $\tilde{s},\tilde{R},\tilde{t}$ indicate the optimized scaling factor, rotation matrix, and translation vector, respectively. $w_{i}$ denotes the confidence weight of 2D keypoint $\mathbf{k}^{i}_{2D}$. We discard any keypoints with confidence below 0.7, as they typically degrade alignment quality. Once the transformation parameters $\tilde{s},\tilde{R},\tilde{t}$ are determined, we apply them to all other SMPL-X sequences under the assumption that they share the same rigid transformation. However, even minor orientation errors can accumulate, leading to physically unrealistic motion trajectories, such as ascending into the sky or descending into the ground. To address this, we adopt GeoCalib [53] to estimate the gravity direction in $W_{env}$, which is then employed to calibrate the SMPL-X to ensure parallel gravity directions, as illustrated in Figure 5(b). For the alignment of Hamer [37], which shares common vertices with the hand parts of SMPL-X, Hamer can also be aligned to $W_{env}$ through the rigid transformation (Equation 4). Additionally, we address the issue of hand occlusion for Hamer by masking occluded hand parts based on the rendered depth from SMPL-X. After the alignments of SMPL-X and Hamer sequences, we place all conditions into $W_{env}$, establishing the global 3D world guidance that allows for rendering concurrently controlled conditions under arbitrary camera trajectories and human motions as shown in Figure 5(c). Finally, these re-rendered conditions are sent to PCDController and RealisDance-DiT for generated outcomes, as shown in Figure 4.
5 Experiments
5.1 Implementation Details
PCDController is trained with the frozen Wan-I2V [54] on multi-resolution images scaled from [$480\times 768$, $512\times 720$, $608\times 608$, $720\times 512$, $768\times 480$] of 81 frames. The learning rate is warmed up to 1e-5 for 400 steps and then fixed. We train the model for 6,000 steps with a batch size of 32, while more training steps would slightly hurt Wan-I2V’s generalization. The training is accomplished with 64 H100 GPUs for 40 hours. We also provided results based on CogVideoX-5B-I2V [64], training for 20k steps with a batch size of 16. During the training, we randomly drop 10% texts, as well as 5% point cloud renderings and Plücker embeddings. For inference, we set the classifier-free guidance scale to 5.0 for textual conditions, keeping other guidance on the default scale 1.0.
Datasets.
To ensure the generalization of PCDController, we collect large-scale training data for camera control, including DL3DV [32], RE10K [74], ACID [34], Co3Dv2 [43], Tartainair [56], Map-Free-Reloc [1], WildRGBD [59], COP3D [49], UCo3D [35]. This comprehensive dataset encompasses a variety of scenarios, featuring both static and dynamic scenes, as well as object-level and scene-level environments. Furthermore, all datasets are annotated with metric-aligned monocular depth through the way proposed in [9] or are provided with ground-truth depth.
5.2 Results of Camera Control
Table 1: Quantitative results of camera control. VBench++ scores (%) are normalized (higher is better). Injected camera features are divided as Plücker ray and point clouds (Pcd). $\dagger$ denotes the results with challenging 360 camera rotations.
[…]
Plücker Pcd Score Consist Consist Quality Quality Flicker Smooth Subject Bg
ViewCrafter [66] $\checkmark$ 85.39 89.69 91.68 55.13 64.33 92.94 97.66 95.59 96.11 0.210 0.117 0.873
SEVA [71] $\checkmark$ 87.39 91.86 93.37 56.79 68.43 95.74 98.59 97.08 97.23 0.077 0.029 0.223
Ours (CogVideoX) $\checkmark$ 86.48 93.17 93.02 55.00 66.75 95.10 98.36 94.45 95.94 0.356 0.162 1.280
Ours (CogVideoX) $\checkmark$ 87.22 91.26 92.44 56.90 69.53 94.79 98.47 96.60 97.79 0.123 0.045 0.346
Ours (Wan-I2V) $\checkmark$ 89.16 94.71 94.93 60.42 72.20 96.51 98.51 97.74 98.29 0.402 0.095 0.728
Ours (Wan-I2V) $\checkmark$ 87.95 91.71 92.97 58.52 71.12 95.51 98.55 97.24 97.96 0.091 0.028 0.211
PCDController $\checkmark$ $\checkmark$ 88.27 92.20 93.37 58.99 71.96 95.56 98.66 97.38 98.01 0.102 0.031 0.246
Ours (Wan-I2V)$\dagger$ $\checkmark$ 1.327 0.551 6.334
[…]
Benchmark.
To evaluate camera control ability, we build an out-of-distribution benchmark with 32 images across various domains, including text-to-image generation [^1] https://github.com/black-forest-labs/flux, real-world [25, 4], object-centric, and human scenes as displayed in Figure 8. Each image features four distinct camera trajectories, resulting in 128 test samples. We used VBench++ [23] to evaluate the video quality, while absolute translation error (ATE), relative translation error (RPE), and relative rotation error (RRE) are used to verify the camera precision. We use VGGT [55] to produce the extrinsic cameras of generated images, which are evaluated with the pre-defined cameras after the trajectory alignment.
Analysis.
We present the quantitative results in Table 1, comparing our model with ViewCrafter [66], SEVA [71], and the CogVideoX [64] version of our framework. The qualitative outcomes are shown in Figure 9. Our experiments demonstrate that point clouds significantly enhance the controllability of both the Wan2.1 and CogVideoX, as verified by the improvement of ATE, RPE, and RRE. Although SEVA achieves precise camera trajectories, it requires massive training with static multi-view data (0.8M iterations), struggling to handle dynamic out-of-distribution scenarios, such as humans and animals, as illustrated in Figure 9. We should clarify that our baseline, Wan-I2V with only Plücker ray, suffers from inferior camera movements. While this setting achieves a strong VBench overall score, it compromises with poor camera metrics. Overall, the proposed PCDContoller achieves the optimal balance between video quality and camera precision. By integrating both Plücker rays and point clouds, it further enhances the performance in challenging scenes featuring substantial viewpoint changes, as validated in Table 1 and Figure 10.
5.3 Results of Unified Camera and Human Motion Control
Table 2: Quantitative results of unified camera and human motion controls. We re-render SMPL-X and Hamer under new camera trajectories for the comparison of RealisDance-DiT [73] without explicitly incorporating point clouds or cameras. $\dagger$ means masking out the foreground point clouds.
[…]
Camera Human Score Consist Consist Quality Quality Flicker Smooth
RealisDance-DiT [73] $\checkmark$ 85.21 93.03 95.34 57.89 68.71 97.44 98.82 0.549 0.195 0.547
PCDController $\checkmark$ 83.19 89.08 91.63 57.23 68.27 95.22 97.71 0.256 0.092 0.661
Uni3C (T2V)$\dagger$ $\checkmark$ $\checkmark$ 83.34 88.45 91.45 57.45 69.84 95.21 97.63 0.296 0.098 1.167
Uni3C (T2V) $\checkmark$ $\checkmark$ 83.16 88.67 91.38 56.79 69.42 95.14 97.57 0.262 0.083 0.606
Uni3C (I2V) $\checkmark$ $\checkmark$ 83.43 89.45 93.05 57.25 67.28 95.70 97.86 0.251 0.093 0.490
Benchmark.
We have developed a new benchmark of 50 videos, each featuring a person performing challenging motions. For guidance, SMPL-X is extracted using GVHMR [48]. Each video is assigned three different types of camera trajectories, resulting in a total of 150 test cases. To ensure that the person remains within the camera’s viewpoint, we employ a follow shooting technique for all test cases, adjusting the camera’s position based on the movement of the SMPL-X center. These noisy and subtle movements further increase the difficulty of camera control. We follow the same camera metrics as mentioned in Section 5.2. To facilitate generalization for motion transfer, RealisDance-DiT is not specifically designed to perfectly recover the first frame aligned with the reference image. Consequently, we remove the metrics that heavily depend on the reference view (I2V subject and background) for fairness.
Analysis.
We show the quantitative results in Table 2, while qualitative results are displayed in Figure 11. To the best of our knowledge, there are currently no publicly available models that effectively address the challenge of unified camera and human motion controls. Therefore, we compare Uni3C against the baseline of RealisDance-DiT [73] and various ablation versions of our model. Notably, while RealisDance-DiT, which focuses solely on human control, achieves the best visual quality, it struggles to produce accurate camera trajectories, resulting in poor camera metrics. In contrast, Uni3C shows good VBench scores alongside impressive camera metrics. Note that the proposed PCDController can also be generalized to the T2V model with comparable quality, featuring the robust generalization of PCDController. Moreover, an interesting insight is revealed from Figure 11 that the aligned SMPL-X characters can further strengthen the camera controllability, while the PCDController trained with I2V formulation also enhances the visual quality and consistency of RealisDance-DiT. This illustrates the complementary features of these two components. Additionally, we show that the proposed Uni3C enables control of detailed hand motions under various camera trajectories as in Figure 12.
5.4 Ablation Study and Exploratory Discussions
Figure 6: Ablation results of Plücker ray and point clouds during the training phase. Point clouds enjoy highly accurate camera control against Plücker ray.
Plücker Ray vs Point Clouds.
As shown in Table 1, point clouds enjoy significantly more precise camera trajectories. We further clarify that video camera control trained with point clouds achieves much faster convergence with lower training loss, as illustrated in Figure 6. Only 1,000 training iterations can hold the general camera trajectory. We empirically find that timing large-scale VDM like Wan-I2V through Plücker ray is difficult. Maybe enabling more trainable parameters would improve the performance, potentially hindering the generalization, which is not considered in this work. Moreover, as verified in the last two rows of Table 1, using both Plücker ray and point clouds improves the results of very challenging camera trajectories.
Point Clouds of Humans.
The point clouds of humans are always frozen in the world space without any motion, which creates a conflict with the human motion conditions provided by SMPL-X. Eliminating this “redundant” information is a straightforward idea to improve motion quality However, as verified in Table 2, while Uni3C-T2V with human-masked point clouds achieves slightly better visual quality, this masking adversely affects camera precision, particularly when humans occupy a significant area of the image. Therefore, retaining the point clouds of humans is essential for effective camera control. Given that Wan2.1 [54] is trained on videos featuring substantial motion, it can generate natural and smooth movements even when the foreground point clouds remain fixed.
Figure 7: Rendering results with and without gravity calibration by GeoCalib [53].
Gravity Calibration.
As mentioned in Section 4.3, gravity calibration is critical for aligning the global 3D world space. Results shown in Figure 7 verify that the calibration can correct the SMPL-X characters aligned with skewed human point clouds and eliminate the error accumulation for humans’ long-distance movements.
Motion Transfer.
We present results of motion transfer achieved by the Uni3C framework in Figure 13. Our model effectively controls both camera trajectories and human motions, even when reference motions are sourced from different videos or distinct domains, such as animation and real-world scenes. Meanwhile, Uni3C can be further extended to generate vivid videos based on other conditions, like text-to-motion guidance or retrieved motions from motion databases. To prove this point, we randomly integrate several motion clips from BABEL [41] and use Uni3C to control both motion and camera, as illustrated in Figure 14. More results are shown on our projected page.
Limitation.
Although Uni3C supports flexible and diverse unified control and motion transfer, it operates under the constraints of predefined camera trajectories and human characters (SMPL-X). Consequently, Uni3C may struggle to produce physically plausible outcomes when human motions conflict with environmental conditions, as illustrated in Figure 15. For instance, if a human’s movement trajectory is blocked by walls, barriers, or other objects, the generated results may exhibit artifacts such as distortion, clipping, or sudden disappearance. This limitation could be mitigated by employing a more advanced human motion generation method that accounts for physical obstructions within the environment.
6 Conclusion
This paper introduced Uni3C, a framework that unifies 3D-enhanced camera and human motion controls for video generation. We first propose the PCDController, demonstrating that lightweight, trainable modules, and rich geometric priors from 3D point clouds can efficiently manage camera trajectories without compromising the inherent capacities of foundational VDMs. This not only enhances generalization but also facilitates versatile downstream applications without joint training. Furthermore, by aligning multi-modal conditions, including both environmental point clouds and human characters in a global 3D world space, we established a coherent framework for jointly controlling camera movements and human animations. Our comprehensive experiments validated the efficacy of Uni3C across diverse datasets, showcasing superior performance in both quantitative and qualitative assessments compared to existing approaches. The significance of our contributions lies not only in improving the state-of-the-art in controllable video generation but also in proposing a robust way to inject multi-modal conditions without the requirements of heavily annotated data. We believe that Uni3C paves the way for more advanced controllable video generation.
Figure 8: Our out-of-distribution benchmark for camera control. The validation set includes generative, human, scene-level, and object-level images with diverse aspect ratios.
Figure 9: Qualitative results of camera control on our benchmark. We compare the proposed PCDController to ViewCrafter [66], SEVA [71], and our model without point cloud guidance. The leftmost image is the reference condition. “full” indicates using both Plücker ray and point clouds as conditions.
Figure 10: Results of the challenging orbital 360 rotations from PCDController. The leftmost images are the reference views.
Figure 11: Results of unified camera and human motion controls. The leftmost images are the reference views, while the first row indicates the aligned 3D world guidance.
Figure 12: Results of unified camera, human motion, and Hamer controls. The leftmost images are the reference views, while the first and second rows indicate the aligned 3D world guidance and Hamer rendering, respectively.
Figure 13: Results of motion transfer. The first row indicates the reference video, while others show our generated videos transferring motions from the reference sequence.
Figure 14: Results transferred from random integrated motion clips of BABEL [41]. The motion sequences are listed on the left, which are executed from light to dark colors.
Figure 15: Failed cases generated by Uni3C. These results are primarily limited by the conflict between human motions and environments.